A boat can travel 5 mile upstream and 5 miles downstream in a total of 6 hours. If the speed of the current is 2 mph, what is the speed of the boat in still water?
Answer by ewatrrr(10682)


Hi
speed of the current is 2 mph
Let (x+2mph) and (x-2mph) represent the speed of the boat down and upstream respectively,
D = r*t OR t = D/r
Question states*** (total time up and down is 6 hr)
5/(x+2) + 5/(x-2) = 6 |multiplying thru by (x+2)(x-2)
5(x+2) + 5(x-2)=6(x+2)(x-2)
10x = 6(x^2 - 4)
6x^2 -10x - 24 = 0
2(3x^2 -5x - 12) = 0
(3x^2 -5x - 12)= 0
factoring
(3x +4)(x-3)= 0 Note:SUM of the inner product(4x) and the outer product(-9x) = -5x
(3x +4)= 0 |tossing out negative solution for speed
(x-3)= 0 x = 3mph, speed of boat in still water
CHECKING our Answer***
5/5 + 5/1 = 6hr